OpenStudy (anonymous):

solve the equation and check solution z^2+6z=-5

OpenStudy (anonymous):

z² + 6z = -5 Add 5. z² + 6x + 5 = 0 Factor. (z + 5)(z + 1) = 0 Zero product property. z = -5, -1 (-5)² + 6(-5) = -5 25 - 30 = -5 (-1)² + 6(-1) = -5 1 - 6 = -5

OpenStudy (lgbasallote):

transpose 5 first... z^2 + 6z+ 5 = 0 factor... (x+5)(x+1) x = -5 x = -1